OpenStudy (anonymous):

Differentiate, y=e ^3x(2x-1), pls help....

OpenStudy (anonymous):

\[e^{3x(2x-1)}\] ?

OpenStudy (anonymous):

so \[(e^{f(x)})' = f'(x)e^{f(x)}\]

OpenStudy (anonymous):

can you do it now ?

OpenStudy (anonymous):

Coolsector......that (2x-1) is not up there with 3x

OpenStudy (anonymous):

if you mean e^(3x)*(2x-1) then the result is e^(3x)*(6x-1) if you meant e^(3x*(2x-1)) then the result is 3*e^(3x*(2x-1))*(4x-1)

OpenStudy (anonymous):

ok .. \[(2x-1)e^{3x}\]

OpenStudy (anonymous):

let \[f(x) = 2x-1 \] \[g(x)=e^{3x}\] \[(f(x)g(x))'=f'(x)g(x)+f(x)g'(x)\]

OpenStudy (anonymous):

can you do it now ?

OpenStudy (anonymous):

Yes thanks, so here you used product rule?

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

thanks a lot .

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

for y=e ^3x(2x-1) it is best to do y=e ^(6x^2-3x) then y'= [e ^(6x^2-3x)](12x-3) or y'= (12x-3)e ^(6x^2-3x) ..... ans......or y'= (12x-3)e ^3x(2x-1 ans...)